OpenStudy (anonymous):

Jerry is trying to round √11 to the nearest tenth. To start, he looked for the perfect squares. 32 and 42 Jerry chose 42 to work with. What did he do wrong?

OpenStudy (anonymous):

3^2 and 4^2

OpenStudy (anonymous):

3^2 = 9 4^2 = 8 9 + 8 = 17 What does that have to do with the original problem?

OpenStudy (anonymous):

Thats what im trying to figure out.

OpenStudy (anonymous):

Well.... 32 divided by 8 = 4

OpenStudy (anonymous):

32 divided by 9 = 3.5 rounded to the nearest tenth = 4

OpenStudy (anonymous):

42 divided by 8 = 5.25 so 5.3 (rounded to the nearest tenth) 42 divided by 9 = 4.6 so 4.7

OpenStudy (anonymous):

That's the closest I can come to the answer. I'm not quite sure what to do next nor how that connects to the answer exactly. I hope I helped you out a little bit.